OpenStudy (anonymous):

Find the perimeter of a rectangle that is 15in x 24in.

OpenStudy (anonymous):

Perimeter= 2(l+b)

jimthompson5910 (jim_thompson5910):

Basically what Champs is saying is that you add the length and width, then you double that result to get the perimeter.

OpenStudy (anonymous):

@jim_thompson5910 39*2?

jimthompson5910 (jim_thompson5910):

yes, which is???

OpenStudy (anonymous):

lets defined : w = width l = lenght P = 2 w + 2 l = 2 * 15 + 2*24 = 30 + 48 =78 in

OpenStudy (anonymous):

78 :}

jimthompson5910 (jim_thompson5910):

you are correct Sabreena7898

OpenStudy (anonymous):

Thanks for your help! @jim_thompson5910

jimthompson5910 (jim_thompson5910):

you're welcome

OpenStudy (anonymous):

Well done sabree..:)

OpenStudy (anonymous):

Shaanks @champ !:}

OpenStudy (anonymous):

Shannks..? :O

OpenStudy (anonymous):

Lmfao, *Thanks{x